OpenStudy (anonymous):

I need help understanding two questions.. Which function has vertex (2, 0) and contains the point (3, 4)? (Points : 1) y = x2 4x + 4 y = 2x2 8x + 8 y = 4x2 4x + 4 y = 4x2 16x + 16

OpenStudy (dtan5457):

I think you missing a bunch of signs on your answers..

OpenStudy (anonymous):

I think its C

OpenStudy (dtan5457):

Please fix your answers or it's impossible to answer lol

OpenStudy (dtan5457):

Yeah. You probably want to either graph this. Or put everything into vertex form. for example y=x^2-4x+4 you don't need to complete the square for this one. it is simply (x-2)^2 vertex=2,0

OpenStudy (dtan5457):

so it's not the first one

OpenStudy (dtan5457):

y=2x^2-8x+8 that's just the first one multiplied by 2, so it's not that either

OpenStudy (dtan5457):

Oh what am I doing..the vertex IS 2,0.

OpenStudy (anonymous):

I used the formula y=a(x-h)^2-k and got a=4 so then I put it into the equation and I got y=4(x-2)^2-0 So I then got y=-8

OpenStudy (dtan5457):

SO it's actually definitely not C, because the vertex is no where near 2,0

OpenStudy (dtan5457):

C's vertex is 0,5,3

OpenStudy (anonymous):

So how do I find the vertex?

OpenStudy (dtan5457):

It's the last one..because y=4x^2-16x+16 y=4(x^2-4x+4) y=4(x-2)^2 vertex=2,0 plug in 3 and you get y=4(3-2)^2 y=4 so D is correct

OpenStudy (anonymous):

Wait but how did you know it was 16??

OpenStudy (dtan5457):

y=4x^2-16x+16 that was the original answer

OpenStudy (dtan5457):

that's what it says on your link

OpenStudy (anonymous):

Yea but do I just go through the answers until I find the right one?

OpenStudy (dtan5457):

Technically, you would put each one into vertex form. Then plug 3 as the x value and see if y comes out as 4

OpenStudy (anonymous):

Okay, can you help with one more problem?

OpenStudy (dtan5457):

sure

OpenStudy (anonymous):

What is the equation of a quadratic function with vertex (0, 0) and containing the points (1, 1), and (1, 1)? (Points : 1) y = x2 y = x2 y = x2 1 y = x2 + 1

OpenStudy (dtan5457):

broken answer choices again

OpenStudy (dtan5457):

containing the points (1, 1), and (1, 1)?

OpenStudy (dtan5457):

It's either x^2 or -x^2, as both vertex is 0,0.

OpenStudy (dtan5457):

Plug in the x coordinates and see if the correct y comes out,

OpenStudy (dtan5457):

Which one do you think it is?

OpenStudy (anonymous):

Idont know I am comletley lost on this one

OpenStudy (dtan5457):

You need to get the points 1,-1 and -1, 1 -(x^2) solve: -(1^2) solve: -(-1^2)

OpenStudy (dtan5457):

*-1,-1

OpenStudy (dtan5457):

Do you get -1 for both answers?

OpenStudy (anonymous):

No just -1,1^2

OpenStudy (dtan5457):

-(-1^2) is not -1?

OpenStudy (anonymous):

Yes it is

OpenStudy (dtan5457):

then your answer is -x^2

OpenStudy (anonymous):

I'm still confused..